OpenStudy (anonymous):

WZ=6, ZX=8, WY=9, find XY

OpenStudy (anonymous):

\[\frac{ WZ }{ YZ } = \frac{ YZ }{ ZX }\]

OpenStudy (anonymous):

And um WZ+ZX=14

OpenStudy (anonymous):

WZ+ZX=WX

OpenStudy (anonymous):

\[\frac{ WZ }{ WY } = \frac{ WY }{ WX }\]

OpenStudy (anonymous):

why you did not use the values

OpenStudy (anonymous):

You have to put the values in .

OpenStudy (anonymous):

\[\frac{ 6 }{ 9 } = \frac{ 9 }{ 14 }\]

OpenStudy (anonymous):

\[\frac{ 8 }{ XY } = \frac{ XY }{ 14 }\]

OpenStudy (anonymous):

I would request you to correct the Question. coz, WY shows = 9 but when you resolve it you are using it as YZ